OpenStudy (anonymous):

Equation of the tangent line to the curve y = 2x ^ 3 +3 x ^ 2 +15 yg parallel to the line 18x - y = 18 is?

myininaya (myininaya):

ok find the slope of 18x-y=18\ assume that one of the tangents of y=2x^3+blah+blah blah is parallel to 18x-y=18 then that means this tangent line has the same slope as the line 18x-y=18 so find y' and set it = to the slope of 18x-y=18 and solve for x

OpenStudy (experimentx):

your curve is quite not clear .. could you please restate the curve and also try using equation

OpenStudy (anonymous):

owh...sory iam new member so I still do not know how to use an open study. :D